# Self-Labeling the Job Shop Scheduling Problem

### v2

@@ Abstract
In this work, we propose a Self-Supervised training strategy specifically designed for combinatorial problems. One of the main obstacles in applying supervised paradigms to such problems is the requirement of expensive target solutions as ground-truth, often produced with costly exact solvers. Inspired by Semi- and Self-Supervised learning, we show that it is possible to easily train generative models by sampling multiple solutions and using the best one according to the problem objective as a pseudo-label. In this way, we iteratively improve the model generation capability by relying only on its self-supervision, completely removing the need for optimality information. We prove the effectiveness of this Self-Labeling strategy on the Job Shop Scheduling (JSP), a complex combinatorial problem that is receiving much attention from the Reinforcement Learning community. We propose a generative model based on the well-known Pointer Network and train it with our strategy. Experiments on popular benchmarks demonstrate the potential of this approach as the resulting models outperform constructive heuristics and current state-of-the-art learning proposals for the JSP.  [^1] The code is available at: https://github.com/AndreaCorsini1/SelfLabelingJobShop
@@ 1 Introduction
The Job Shop Problem is a classic optimization problem with many practical applications in industry and academia [51]. At its core, the JSP involves scheduling a set of jobs onto a set of machines, where each job has to be processed on the machines following a specific order. The goal of this problem is generally to minimize the makespan, i.e., the time required to complete all jobs [39].
@@
Over the years, various approaches have been developed to tackle the JSP. A common one is adopting exact methods such as Mixed Integer Programming solvers (MIP). However, these methods often fail to provide either optimal or high-quality solutions on medium and large instances [29]. As a remedy, metaheuristics have been extensively investigated [35, 13, 24], and constitute an alternative to exact methods. Although state-of-the-art metaheuristics like [36] can provide quality solutions in a fair amount of time, typically minutes, they are rather complex to implement and is hard to reproduce their results [6]. Therefore, due to their lower complexity, and under tighter time constraints, Priority Dispatching Rules (PDR) [39] are often preferred, despite their lower quality. The main issue of PDRs is their tendency to perform well in some cases and poorly in others, primarily due to their blind prioritization schema based on rigid hardcoded and hand-crafted rules [21].
@@
Therefore, recent works have started to investigate Machine Learning (ML) to enhance or replace some of these approaches, primarily focusing on PDRs. ML-based approaches, specifically Reinforcement Learning (RL) ones, demonstrated to deliver higher-quality solutions than classic PDRs at the cost of a small increase in execution time [32]. Despite the potential of RL [49, 25], training RL agents is a complex task [42, 33], is sensitive to hyper-parameters [40], and has reproducibility issues [22].
@@
Contrary, supervised learning is less affected by these issues but heavily relies on expensive annotations [31]. This is particularly problematic in combinatorial problems, where annotations in the form of (near-) optimal information are generally produced with expensive (exact) solvers. To contrast labeling cost and improve generalization, Semi-Supervised [45] and Self-Supervised [31] are gaining popularity due to their ability to learn from unlabeled data. Despite this premise, little to no application of these paradigms can be found in the JSP and the combinatorial literature [5, 9].
@@
Motivated by these observations, we introduce a novel self-supervised training strategy that is less complex than most RL algorithms, does not require the formulation of the Markov Decision Process, and relies only on model-generated information, thus removing the need for expensive annotations. Our proposal is based on two assumptions: i) we suppose to be able to generate multiple solutions for a problem, a common characteristic of generative models [37]; and ii) we suppose it is possible to discriminate solutions based on the problem objective. When these assumptions are met, we train a model by generating multiple solutions and using the best one according to the problem objective as a pseudo-label [30]. This procedure borrows from Semi-Supervised the idea of pseudo-labeling, but does not require any external and expansive annotation as in Self-Supervised learning. Hence, we refer to it as a Self-Labeling training strategy.
@@
We demonstrate the viability of our training procedure in the context of the JSP. Similar to PDRs, we cast the generation of solutions as a sequence of decisions, where at each decision one operation is scheduled in the solution under construction. This is achieved with a popular architecture known as Pointer Network (PN) for dealing with sequences of decisions in combinatorial problems [47, 27, 49]. We train our PN on 30 000 instances of different sizes by generating multiple parallel solutions and using the one with the minimum makespan to update the model. This training strategy produces models outperforming state-of-the-art learning proposals and other conventional JSP algorithms on two popular benchmark sets. Additionally, our methodology proved to be robust and effective in various training regimes. Summarizing, the contributions of this work are:
  • We propose a novel Self-Labeling training strategy. Due to its assumptions, this effective strategy for the JSP can be applied as-is to other combinatorial problems.
  • We present a generative model based on the Pointer Network which produces high-quality JSP solutions in a matter of seconds when trained with our strategy.
@@
The remainder of this work is organized as follows: Section 2 reviews related ML works; Section 3 formalizes the JSP; Section 4 introduces the proposed PN and training strategy; Section 5 compares our trained PN with other approaches; Section 6 studies aspects of our proposal; and Section 7 and 8 close with general considerations, some limitations, and potential future directions.
@@ 2 Related Works
Motivated by the success of supervised learning, various studies have employed exact methods to generate optimality information for synthetic JSP instances, enabling the learning of valuable scheduling patterns. For example, [26] used imitation learning [5] to derive superior dispatching rules from optimal solutions, highlighting the limitations of learning solely from optimal solutions. [28] proposed a joint utilization of optimal solutions and Lagrangian terms during training to better capture JSP constraints. Whereas, in [14], a Recurrent Neural Network (RNN) learns to predict the quality of machine permutations, generated during training with a MIP, for guiding a metaheuristic. Despite their quality, these proposals are limited by their dependence on costly optimality information.
@@
To avoid the need for optimality, other works have turned to reinforcement learning. The advantage of RL lies in its independence from costly optimal information, requiring only the correct formulation of the Markov Decision Process [42]. Several works focused on creating better neural PDRs, showcasing the effectiveness of policy-based methods [49, 38, 25, 11]. All these methods adopted a Proximal Policy Optimization algorithm [40] and proposed different architectures or training variations. For instance, [49, 38] adopted Graph Neural Networks (GNN), [11] proposed a rather complex Transformer [46], while [25] used RNNs and curriculum learning [53]. Differently, [20] presented a Double Deep Q-Network approach, proving the applicability of value-based methods.
@@
Other works employed RL to enhance other algorithmic approaches. In [12], an iterative RL-based approach was proposed to rewrite regions of JSP solutions selected by a learned policy. Whereas [17] utilized a Deep Q-learning approach to control three points of interventions within metaheuristics, while [50] recently presented a RL-guided improvement heuristic. Differently, [44] presented a hybrid imitation learning and policy gradient approach coupled with Constraint Programming (CP) for outperforming PDRs and a CP solver.
3 Job Shop
@@
In the Job Shop Problem, we are given a set of jobs $J=\{J_{1},\dots,J_{n}\}$, a set of machines $M=\{M_{1},\dots,M_{m}\}$, and a set of operations $O=\{1,\dots,o\}$. Each job $j\in J$ comprises a sequence of $m_{j}$ consecutive operations $O_{j}=(l_{j},\,\dots,l_{j}+m_{j})\subset O$ indicating the order in which the job must be performed on machines, where $l_{j}$ is the index of the first operation of the job ($l_{j}=1+\sum_{i=1}^{j-1}m_{i}$). An operation $i\in O$ has to be performed on machine $\mu_{i}\in M$ for an uninterrupted amount of time $\tau_{i}\in\mathbb{R}_{\geq 0}$. Additionally, machines can handle one operation at a time. The objective of the JSP is to provide an order to operations on each machine, such that the precedences within operations are respected and the time required to complete all jobs (makespan) is minimized. Formally, we indicate with $\pi$ a JSP solution comprising a permutation of operations for each machine and use $C_{i}(\pi)$ to identify the completion time of operation $i\in O$ in $\pi$.
@@
A JSP instance can be represented with a disjunctive graph $G=(V,A,E)$ [3], where: $V$ contains one vertex for each operation $i\in O$; $A$ is the set of conjunctive (directed) arcs connecting consecutive operations of jobs (these arcs reflect the order in which operations must be performed); $E$ is the set of disjunctive (undirected) edges connecting operations to perform on the same machines. When the JSP is represented as a disjunctive graph $G$, finding a solution means providing a direction to all the edges in $E$, such that the resulting graph is directed and acyclic (all precedences are respected). As in related works [49, 38], we augment the disjunctive graph by associating to each vertex a set of 15 features $x_{i}\in\mathbb{R}^{15}$ describing information about operation $i\in O$. For lack of space, we present these features in Appendix A.
@@
Figure 1: The sequences of decisions for constructing solutions in a JSP instance with two jobs ($J_{1}$ and $J_{2}$) and two machines (identified with different colors).
3.1 Constructing JSP Solutions
@@
Feasible JSP solutions can be constructed step-by-step as a sequence of decisions, a common approach adopted by PDRs [39] and RL algorithms [49, 25]. The construction of solutions can be visualized using a branch-decision tree, shown at the bottom of Figure 1. Each path from the root node ($R$) to a leaf node presents a particular sequence of $o=|O|$ decisions and leads to a valid JSP solution, such as the one highlighted in gray. At every node in the tree, one uncompleted job needs to be selected, and its ready operation is scheduled in the partial solution being constructed. Due to the constraints imposed by the jobs (i.e., the conjunctive arcs) and the partial solution $\pi_{t}$ constructed up to decision $t$, there is at most one operation $o(t,\,j)\in O_{j}$ that is ready to be scheduled for any job $j\in J$. Thus, by selecting a job $j$, we uniquely identify an operation $o(t,\,j)$ that will be scheduled by appending it to the partial permutation of its machine. It is important to note that once a job is completed, it cannot be selected again, and such a situation is identified with a cross in Figure 1.
We also highlight that there exist multiple paths in the branch-decision tree that lead to the same solution. This indicates that, within the context of the JSP, certain solutions are more probable than others. In some instances, due to the distribution of processing times $\tau_{i}$, near-optimal solutions may be found among the less likely paths, making the construction of quality solutions a more complex task. This explains why it is more difficult for rigid constructive heuristics like standard PDRs to create high-quality solutions when such solutions result from very unlikely paths.
4 Pointer Network & Self-Labeling Strategy
Herein, we present the Pointer Network used to tackle the JSP and the proposed self-supervised training strategy.
4.1 Proposed Pointer Network
@@
As described in Section 3.1, we tackle the JSP as a sequence of decisions. Thus, we propose a Pointer Network [47] whose goal is to select the right job at each decision step $t$. Recall that the PN is an encoder-decoder architecture to generate sequences of decisions in combinatorial problems [34, 27]. Formally, our PN learns a function $f_{\theta}(\cdot)$, parametrized by $\theta$, that estimates the probability $p_{\theta}(\pi|G)$ of a solution $\pi$ of being of high-quality as a product of probabilities:
$p_{\theta}(\pi|G)=\prod_{t=1}^{o}f_{\theta}(j\,|\,\pi_{t},G),$ (1)
@@
where $f_{\theta}(j\,|\pi_{t},G)$ gives the probability of selecting job $j\in J$ conditioned on the instance $G$ and the partial solution $\pi_{t}$ at step $t$. By learning $f_{\theta}$, it is possible to construct solutions autoregressively with a high chance of being of high quality.
@@
Our PN works by taking in input JSP instances represented as a disjunctive graph $G$, and produces single deterministic or multiple randomized solutions depending on the adopted construction strategy. The encoder of our PN creates with a single forward computation of $G$ an embedded representation for all the operations in $O$. Whereas the decoder uses the embeddings of operations and the partial solution $\pi_{t}$ to produce a probability for each job of being selected at step $t$. After selecting a job $j$, its ready operation $o(t,\,j)$ is scheduled in $\pi_{t}$ to produce the partial solution for step $t+1$. In the following, we present the details of both the encoder and decoder.
Encoder.
@@
Its role is to encode meaningful patterns within an instance into the embedded representation of operations. The encoder can be embodied by any architecture, like a Feedforward Neural Network (FNN), that transforms the feature of operations $x_{i}\in\mathbb{R}^{15}$ (see Appendix A for the features) into embeddings $e_{i}\in\mathbb{R}^{h}$, without necessarily accounting for information contained in the graph $G$. As in related works [49, 38], we also encode the relationships among operations present in the disjunctive graph. Thus, we equip the encoder with Graph Neural Networks [48], which enables the embeddings to incorporate information related to the topological structure of $G$. In our encoder, we stack two layers of Graph Attention Network [8] (GAT) as follows:
@@
Table 1: The features of a context vector $c_{j}\in\mathbb{R}^{11}$ that describes the status of a job $j$ within a partial solution $\pi_{t}$. Recall that $o(t,j)$ is the ready operation of job $j$ at step $t$, $\mu_{o(t,\,j)}$ its machine, and $o(t,j)-1$ its predecessor.
@@
1 $C_{o(t,\,j)-1}(\pi_{t})$ minus the completion time of machine $\mu_{o(t,\,j)}$. The idle time of job $j$ if scheduled on its next machine $\mu_{o(t,\,j)}$ at time $t$. Negative if the job could have started earlier.
2 $C_{o(t,\,j)-1}(\pi_{t})$ divided by the makespan of $\pi_{t}$. How close job $j$ is to the makespan of the partial solution $\pi_{t}$.
3 $C_{o(t,\,j)-1}(\pi_{t})$ minus the average completion time of all jobs. How early/late job $j$ completes w.r.t. the average among all jobs in $\pi_{t}$.
4 - 6 The difference between $C_{o(t,\,j)-1}(\pi_{t})$ and the $1^{st}$, $2^{nd}$, and $3^{rd}$ quartile computed among the completion time of all jobs. Describe the relative completion of $j$ w.r.t. other jobs. These features complement that with ID = $3$.
7 The completion time of machine $\mu_{o(t,\,j)}$ divided by the makespan of the partial solution $\pi_{t}$. How close the completion time of machine $\mu_{o(t,\,j)}$ is to the makespan of the partial solution $\pi_{t}$.
8 The completion time of machine $\mu_{o(t,\,j)}$ minus the average completion of all machines in $\pi_{t}$. How early/late machine $\mu_{o(t,\,j)}$ completes w.r.t. the average among machines in $\pi_{t}$.
@@
Decoder.
Its role is to produce at any step $t$ the probability of selecting each job with two logically distinct components:
  • Memory Network: generates a state $s_{j}\in\mathbb{R}^{d}$ for each job $j\in J$ from the partial solution $\pi_{t}$. This is achieved by first extracting from $\pi_{t}$ a context vector $c_{j}$ (similarly to [27]), which contains eleven hand-crafted features providing useful cues about job $j$ in the partial solution. Note that these features are an integral part of our architecture and we refer to Table 1 for their definition and meaning. Then, these vectors are fed into a Multi-Head Attention layer (MHA) [46] followed by a non-linear projection to produce jobs’ states:
@@
    Then, these scores $z_{j}\in\mathbb{R}$ are transformed into probabilities with a Softmax function: $p_{j}=e^{z_{j}}\,/\,\sum\nolimits_{b\in J}e^{z_{b}}$. Finally, the decision of which job to select at $t$ is made with a sampling strategy, as explained next.
Sampling solutions.
@@
To generate solutions, we employ a probabilistic approach for deciding the job selected at any step. Specifically, we randomly sample a job $j$ with a probability proportional to $p_{j}$, which is produced at step $t$ by our decoder. We also prevent the selection of completed jobs by setting their probabilities to zero before sampling, i.e., we set $p_{j}=0$ for completed $j$. Note how this probabilistic selection is autoregressive, meaning that sampling a job depends on $p_{j}$ which is a function of $e_{o(t,\,j)}$ and $s_{j}$ resulting from earlier decisions (see Eq. 3 and 4). There exist various strategies for sampling from autoregressive models, including top-k [23], nucleus [23], and random sampling [4]. After preliminary analysis, we found that these strategies exhibit similar effectiveness in our setting. Therefore, we adopt the simplest strategy for training and testing, which is random sampling, i.e., sampling a job $j$ with probability $p_{j}$ at random. Lastly, we use our PN to generate $\beta$ parallel solutions for an instance. This is achieved by keeping $\beta$ separate partial solutions and sampling at any step a job for each one independently.
4.2 Self-Labeling Training Strategy

### v3

@@ Abstract
This work proposes a self-supervised training strategy designed for combinatorial problems. An obstacle in applying supervised paradigms to such problems is the need for costly target solutions often produced with exact solvers. Inspired by semi- and self-supervised learning, we show that generative models can be trained by sampling multiple solutions and using the best one according to the problem objective as a pseudo-label. In this way, we iteratively improve the model generation capability by relying only on its self-supervision, eliminating the need for optimality information. We validate this Self-Labeling Improvement Method (SLIM) on the Job Shop Scheduling (JSP), a complex combinatorial problem that is receiving much attention from the neural combinatorial community. We propose a generative model based on the well-known Pointer Network and train it with SLIM. Experiments on popular benchmarks demonstrate the potential of this approach as the resulting models outperform constructive heuristics and state-of-the-art learning proposals for the JSP. Lastly, we prove the robustness of SLIM to various parameters and its generality by applying it to the Traveling Salesman Problem.
@@ 1 Introduction
The Job Shop Problem (JSP) is a classic optimization problem with many practical applications in industry and academia [56]. At its core, the JSP entails scheduling a set of jobs across multiple machines, where each job has to be processed on the machines following a specific order. The goal of this problem is generally to minimize the makespan: the time required to complete all jobs [40].
@@
Over the years, various approaches have been developed to tackle the JSP. A common one is adopting exact methods such as Mixed Integer Programming solvers (MIP). However, these methods often struggle to provide optimal or high-quality solutions on medium and large instances in a reasonable timeframe [29]. As a remedy, metaheuristics have been extensively investigated [36, 13, 24] and constitute an alternative to exact methods. Although state-of-the-art metaheuristics like [37] can rapidly provide quality solutions, typically within minutes, they are rather complex to implement and their results can be difficult to reproduce [6]. Therefore, due to their lower complexity, Priority Dispatching Rules (PDR) [40] are frequently preferred in practical applications with tighter time constraints. The main issue of PDRs is their tendency to perform well in some cases and poorly in others, primarily due to their rigid, hand-crafted prioritization schema based on hardcoded rules [21].
@@
Recent works have increasingly investigated Machine Learning (ML) to enhance or replace some of these approaches, primarily focusing on PDRs. ML-based approaches, specifically Reinforcement Learning (RL) ones, proved to deliver higher-quality solutions than classic PDRs at the cost of a small increase in execution time [33]. Despite the potential of RL [54, 25], training RL agents is a complex task [45, 34], is sensitive to hyper-parameters [43], and has reproducibility issues [22].
@@
Contrary, supervised learning is less affected by these issues but heavily relies on expensive annotations [32]. This is particularly problematic in combinatorial problems, where annotations in the form of (near-) optimal information are generally produced with expensive exact solvers. To contrast labeling cost and improve generalization, semi-supervised [48] and self-supervised [32] are gaining popularity in many fields due to their ability to learn from unlabeled data. Despite this premise, little to no application of these paradigms can be found in the JSP and the combinatorial literature [5, 9].
@@
Motivated by these observations, we introduce a novel self-supervised training strategy that is simpler than most RL algorithms, does not require the formulation of the Markov Decision Process [45], and relies only on model-generated information, thus removing the need for expensive annotations. Our proposal is based on two weak assumptions: i) we suppose to be able to generate multiple solutions for a problem, a common characteristic of generative models [38]; and ii) we suppose it is possible to discriminate solutions based on the problem objective. When these assumptions are met, we train a model by generating multiple solutions and using the best one according to the problem objective as a pseudo-label [31]. This procedure draws on the concept of pseudo-labeling from semi-supervised learning, but it does not require any external and expansive annotation as in self-supervised learning. Hence, we refer to it as a SLIM: Self-Labeling Improvement Method.
We prove the effectiveness of SLIM primarily within the context of Job Shop, a challenging scheduling problem with many baseline algorithms [1, 36] and established benchmarks [44, 46, 16]. As recognized in other works [54, 39, 25], focusing on the JSP is crucial because its study helps address related variants, such as Dynamic JSP [51] and Flow Shop [42], while also establishing a concrete base for tackling more complicated scheduling problems like Flexible JSP [53].
@@
Similar to PDRs, we cast the generation of solutions as a sequence of decisions, where at each decision one job operation is scheduled in the solution under construction. This is achieved with a generative model inspired by the Pointer Network [50], a well-known architecture for dealing with sequences of decisions in combinatorial problems [5]. We train our model on random instances of different sizes by generating multiple parallel solutions and using the one with the minimum makespan to update the model. This training strategy produces models outperforming state-of-the-art learning proposals and other conventional JSP algorithms on popular benchmark sets. Furthermore, our methodology demonstrates robustness and effectiveness across various training regimes, including a brief analysis on the Traveling Salesman Problem to emphasize the broader applicability of SLIM beyond scheduling. In summary, the contributions of this work are:
  • Our key contribution is the introduction of SLIM, a novel self-labeling improvement method for training generative models. Thanks to its minimal assumptions, SLIM can be easily applied as-is to other combinatorial problems.
  • Additionally, we present a generative encoder-decoder architecture capable of generating high-quality solutions for JSP instances in a matter of seconds.
@@
The remainder of this work is organized as follows: Sec. 2 reviews related ML works; Sec. 3 formalizes the JSP; Sec. 4 introduces our generative model and SLIM; Sec. 5 compares our approach with others; Sec. 6 studies additional aspects of our proposal; and Secs. 7 and 8 close with general considerations, some limitations, and potential future directions.
@@ 2 Related Works
Motivated by the success of supervised learning, various studies have employed exact methods to generate optimality information for synthetic JSP instances, enabling the learning of valuable scheduling patterns. For example, [26] used imitation learning [5] to derive superior dispatching rules from optimal solutions, highlighting the limitations of learning solely from optimal solutions. [28] proposed a joint utilization of optimal solutions and Lagrangian terms during training to better capture JSP constraints. Whereas, in [14], the authors trained a Recurrent Neural Network (RNN) to predict the quality of machine permutations – generated during training with an MIP – for guiding a metaheuristic. Despite their quality, these proposals are limited by their dependence on costly optimality information.
@@
To avoid the need for optimality, other works have turned to reinforcement learning. The advantage of RL lies in its independence from costly optimal information, requiring only the effective formulation of the Markov Decision Process [45]. Several works focused on creating better neural PDRs, showcasing the effectiveness of policy-based methods [54, 39, 25, 11]. All these methods adopted a Proximal Policy Optimization algorithm [43] and proposed different architectures or training variations. For instance, [54, 39] adopted Graph Neural Networks (GNN), [11] proposed a rather complex Transformer [49], while [25] used RNNs and curriculum learning [58]. Differently, [20] presented a Double Deep Q-Network approach, proving the applicability of value-based methods.
@@
Other works employed RL to enhance other algorithmic approaches. In [12], an iterative RL-based approach was proposed to rewrite regions of JSP solutions selected by a learned policy. Whereas [17] utilized a Deep Q-learning approach to control three points of interventions within metaheuristics, and [55] recently presented an RL-guided improvement heuristic. Differently, [47] presented a hybrid imitation learning and policy gradient approach coupled with Constraint Programming (CP) for outperforming PDRs and a CP solver.
3 Job Shop Scheduling
@@
In the Job Shop Problem, we are given a set of jobs $J=\{J_{1},\dots,J_{n}\}$, a set of machines $M=\{M_{1},\dots,M_{m}\}$, and a set of operations $O=\{1,\dots,o\}$. Each job $j\in J$ comprises a sequence of $m_{j}$ consecutive operations $O_{j}=(l_{j},\,\dots,l_{j}+m_{j}-1)\subset O$ indicating the order in which the job must be performed on machines, where $l_{j}=1+\sum_{i=1}^{j-1}m_{i}$ is the index of its first operation (e.g., $l_{1}=1$ for $J_{1}$ and $l_{2}=3$ for $J_{2}$ in Fig. 1). An operation $i\in O$ has to be performed on machine $\mu_{i}\in M$ for an uninterrupted amount of time $\tau_{i}\in\mathbb{R}_{\geq 0}$. Additionally, machines can handle one operation at a time. The objective of the JSP is to provide an order to operations on each machine, such that the precedences within operations are respected and the time required to complete all jobs (makespan) is minimized. Formally, we indicate with $\pi$ a JSP solution comprising a permutation of operations for each machine and use $C_{i}(\pi)$ to identify the completion time of operation $i\in O$ in $\pi$.
@@
A JSP instance can be represented with a disjunctive graph $G=(V,A,E)$ [3], where: $V$ contains one vertex for each operation $i\in O$; $A$ is the set of directed arcs connecting consecutive operations of jobs reflecting the order in which operations must be performed; $E$ is the set of disjunctive (undirected) edges connecting operations to perform on the same machines. When the JSP is represented as a disjunctive graph $G$, finding a solution means providing a direction to all the edges in $E$, such that the resulting graph is directed and acyclic (all precedences are respected). As in related works [54, 39], we augment the disjunctive graph by associating to each vertex a set of $15$ features $x_{i}\in\mathbb{R}^{15}$ describing information about operation $i\in O$. For lack of space, we present these features in App. A.
@@
Figure 1: The sequences of decisions for constructing solutions in a JSP instance with two jobs ($J_{1}$ and $J_{2}$) and two machines (identified in green and red). Best viewed in colors.
3.1 Constructing Feasible Job Shop Solutions
@@
Feasible JSP solutions can be constructed step-by-step as a sequence of decisions, a common approach adopted by PDRs [40] and RL algorithms [54, 25]. The construction of solutions can be visualized using a branch-decision tree, shown at the bottom of Fig. 1. Each path from the root node ($R$) to a leaf node presents a particular sequence of $o=|O|$ decisions and leads to a valid JSP solution, such as the one highlighted in gray. At every node in the tree, one uncompleted job needs to be selected, and its ready operation is scheduled in the solution being constructed. Due to the precedence constraints and the partial solution $\pi_{<t}$ constructed up to decision $t$, there is at most one operation $o(t,\,j)\in O_{j}$ that is ready to be scheduled for any job $j\in J$. Thus, by selecting a job $j$, we uniquely identify an operation $o(t,\,j)$ that will be scheduled by appending it to the partial permutation of its machine. Notice that once a job is completed, it cannot be selected again, and such a situation is identified with a cross in Fig. 1.
We emphasize that diverse paths can lead to the same solution, indicating certain solutions are more frequent than others. In some instances, near-optimal solutions may lie at the end of unlikely and hardly discernible paths, making their construction more challenging. This could explain why PDRs, which rely on rigid, predefined rules, sometimes struggle to generate quality solutions, whereas neural algorithms – able to leverage instance- and solution-wide features – may perform better.
4 Methodology: Generative Model & Self-Labeling
@@
As described in Sec. 3.1, we tackle the JSP as a sequence of decisions. Thus, we propose a generative model inspired by the Pointer Network [50] (a well-known encoder-decoder architecture to generate sequences of decisions) whose goal is to select the right job at each decision step $t$. Formally, our model learns a function $f_{\theta}(\cdot)$, parametrized by $\theta$, that estimates the probability $p_{\theta}(\pi|G)$ of a solution $\pi$ being of high-quality, expressed as a product of probabilities:
$p_{\theta}(\pi|G)=\prod_{t=1}^{|O|}f_{\theta}(y_{t}\,|\,\pi_{<t},G),$ (1)
@@
where $f_{\theta}(y_{t}\,|\pi_{<t},G)$ gives the probability of selecting job $y_{t}\in J$ for creating $\pi$, conditioned on the instance $G$ and the partial solution $\pi_{<t}$ constructed up to step $t$. By accurately learning $f_{\theta}(\cdot)$, the model can construct high-quality solutions autoregressively. The following sections present the proposed encoder-decoder architecture and the self-labeling strategy to train this generative model.
4.1 The Generative Encoder-Decoder Architecture
@@
Overview. Our model processes JSP instances represented as a disjunctive graph $G$ and produces single deterministic or multiple randomized solutions depending on the adopted construction strategy. The encoder operates at the instance level, creating embedded representations for all the operations in $O$ with a single forward computation of $G$ (hence, no autoregression). Whereas the decoder operates at the solution level (job and machine level), using the embeddings of ready operations and the partial solution $\pi_{<t}$ to produce a probability of selecting each job at step $t$. Recall that after selecting a job $j$, its ready operation $o(t,\,j)$ is scheduled in $\pi_{<t}$ to produce the partial solution for step $t+1$.
@@
Encoder. It captures instance-wide relationships into the embeddings of operations, providing the decoder with a high-level view of the instance characteristics. The encoder can be embodied by any architecture, like a Feedforward Neural Network (FNN), that transforms raw operation features $x_{i}\in\mathbb{R}^{15}$ (see App. A for the features) into embeddings $e_{i}\in\mathbb{R}^{h}$. As in related works [54, 39], we also encode the relationships among operations present in the disjunctive graph. We thus equip the encoder with Graph Neural Networks [52], allowing the embeddings to incorporate topological information of $G$. In our encoder, we stack two Graph Attention Network (GAT) layers [8] as follows:
@@
Table 1: The features of a context vector $c_{j}\in\mathbb{R}^{11}$ that describes the status of a job $j$ within a partial solution $\pi_{<t}$. Recall that $o(t,j)$ is the ready operation of job $j$ at step $t$, $\mu_{o(t,\,j)}$ its machine, and $o(t,j)-1$ its predecessor.
@@
1 $C_{o(t,\,j)-1}(\pi_{<t})$ minus the completion time of machine $\mu_{o(t,\,j)}$. The idle time of job $j$ if scheduled on its next machine $\mu_{o(t,\,j)}$ at time $t$. Negative if the job could have started earlier.
2 $C_{o(t,\,j)-1}(\pi_{<t})$ divided by the makespan of $\pi_{<t}$. How close job $j$ is to the makespan of the partial solution $\pi_{<t}$.
3 $C_{o(t,\,j)-1}(\pi_{<t})$ minus the average completion time of all jobs. How early/late job $j$ completes w.r.t. the average among all jobs in $\pi_{<t}$.
4 - 6 The difference between $C_{o(t,\,j)-1}(\pi_{<t})$ and the $1^{st}$, $2^{nd}$, and $3^{rd}$ quartile computed among the completion time of all jobs. Describe the relative completion of $j$ w.r.t. other jobs. These features complement that with ID = $3$.
7 The completion time of machine $\mu_{o(t,\,j)}$ divided by the makespan of the partial solution $\pi_{<t}$. How close the completion time of machine $\mu_{o(t,\,j)}$ is to the makespan of the partial solution $\pi_{<t}$.
8 The completion time of machine $\mu_{o(t,\,j)}$ minus the average completion of all machines in $\pi_{<t}$. How early/late machine $\mu_{o(t,\,j)}$ completes w.r.t. the average among machines in $\pi_{<t}$.
@@
Decoder. It produces at any step $t$ the probability of selecting each job from the embeddings $e_{i}$ and solution-related features. The encoder is logically divided into two distinct components:
  • Memory Network: generates a state $s_{j}\in\mathbb{R}^{d}$ for each job $j\in J$ from the partial solution $\pi_{<t}$. This is achieved by first extracting from $\pi_{<t}$ a context vector $c_{j}$ (similarly to [27]), which contains eleven hand-crafted features providing useful cues about the status of job $j$. We refer to Tab. 1 for the definition and meaning of these features. Then, these vectors are fed into a Multi-Head Attention layer (MHA) [49] followed by a non-linear projection to produce jobs’ states:
@@
    Then, these scores $z_{j}\in\mathbb{R}$ are transformed into softmax probabilities: $p_{j}=e^{z_{j}}\,/\,\sum\nolimits_{b\in J}e^{z_{b}}$. The final decision on which job to select at $t$ is made using a sampling strategy, as explained next.
@@
Sampling solutions. To generate solutions, we employ a probabilistic approach for deciding the job selected at any step. Specifically, we randomly sample a job $j$ with a probability $p_{j}$, produced at step $t$ by our decoder. We also prevent the selection of completed jobs by setting their scores $z_{j}=-\infty$ to force $p_{j}=0$ before sampling. Note how this probabilistic selection is autoregressive, meaning that sampling a job depends on $p_{j}$ which is a function of $e_{o(t,\,j)}$ and $s_{j}$ resulting from earlier decisions (see Eqs. 3 and 4). Various strategies exist for sampling from autoregressive models, including top-k [23], nucleus [23], and random sampling [4]. Preliminary experiments revealed that these strategies perform similarly, with average optimality gap variations within $0.2\%$ on benchmark instances. However, top-k and nucleus sampling introduce brittle hyperparameters that can hinder training convergence if not managed carefully. Therefore, we adopt the simplest strategy for training and testing, which is random sampling: sampling a job $j$ with probability $p_{j}$ at random. Note however that the sampling strategy remains a flexible design choice, not inherently tied to our methodology.
4.2 SLIM: Self-Labeling Improvement Method